The sum of the ages of Joe and Jim is 92. Eight years from now, Joe will be twice Jim's age. Find Joe's present age?
Answer by JimboP1977(311)

Joe's present age = X
Jim's present age = Y
X+Y=92 (1)
X+8 = 2*(Y+8) (2)
X+8 = 2Y+16 (2)
X = 2Y + 8 (2)
Substitute (2) into (1)
2Y+8+Y = 92
3Y+8 =92
3Y = 84
Y = 28
X =92-28
X = 64
Does this make any sense? Please get back to me.